Reports 1-1 of 1 Clear search Modify search
OBS (General)
keiko.kokeyama - 20:05 Wednesday 08 April 2020 (14048) Print this report
Comment to Shift Summary - Day Apr. 8, 2020 (14046)

Lucia, Keiko, Ushiba

Summary of the recovery work:

Situation after lockloss: GR X and GR Y TRANS were very low (0.3~0.5) possibly locking with higher order modes whereas IR arm was good.

[Initial Alignemnt]
X arm:
First I tried to change the GR beam to align the IR. POP POM pico was moved especially in YAW (~ +-100 steps). However GR TRANS coudn't be maxmized to 1. With suggestions from Nakano-kun, after all PR3, PR2 and IMMT1 were moved to align X arm GR and IR. Now I think maybe IMMT1 could be moved but hasn't been moved. X arm ADS was failing, but after the alignment it worked.

Y arm:
Y arm GR and IR locked but the ADS didn't work in the beginning. ETM COMM DIFF dither CLK GAIN were on and they may have been messing up the Y arm ADS.. Also, I did a burt of the ADS model to a wrong timing, the Y arm dither signal was using AS DC instead of TMSY signal. That was also messing up the EY ADS. It took some time to notice them and fix them.

When Y arm initial alignment, after ALS and IR length lock, we manually centered the OMC QPDs with OMMMT/OSTM pico and OPTIC ALIGN sliders (these QPD SUMs have to be ~2900 and ~2500 for QPDV1 and QPDV2, respectively). OMC was locked OK, however, when ENGAGE_OMC_ADS, there were strange kicks and jumps of OMMT2 and OSTM when we engage OMC_LSC_LOCKED. We repeated the procedures and saw the kicks for 2-3 times. However after some trials, it didn't happen again. We left as it was and we don't know what were the problems.

IFO recovered at around 8pm.

Search Help